# The compact object of HESS J1731-347 and its implication on neutron star matter

### v2

@@ I Introduction
Recent observation of a compact object inside the supernova remnant HESS J1731-347 has provided an estimation of its mass and radius to be $M=0.77^{+0.20}_{-0.17}M_{\odot}$ and $R=10.4^{+0.86}_{-0.78}$ km at $1\sigma$ credible levels, respectively Doroshenko et al. (2022). Understandably, this observation has generated a lot of attention from the nuclear physics community working on the dense matter physics beyond the normal nuclear matter density Brodie and Haber (2023); Huang et al. (2024); Li and Sedrakian (2023); Kubis et al. (2023); Koehn et al. (2025); Miao et al. (2024). Different possibilities have been explored to characterize this object, such as, a strange quark or a hybrid star Di Clemente et al. (2024); Horvath et al. (2023); Oikonomou and Moustakidis (2023); Rather et al. (2023); Yuan and Li (2024); Laskos-Patkos et al. (2024); Gao et al. (2024); Mariani et al. (2024), or a dark matter admixed compact star Sagun et al. (2023); Routaray et al. (2023). A typical neutron star (NS) usually has a mass within the range of $\sim 1-2M_{\odot}$. NSs being extremely dense compact objects have profound effects on our understanding of dense nuclear matter higher than nuclear saturation density ($\rho_{0}$) Glendenning (1997). The equation of state (EOS) of highly dense matter is connected to the observations of the masses and radii of the compact objects through the solutions of the TOV equations. Hence, each new source with observed mass and radius allows us to test and constrain our models of nuclear physics and increase precision over the previous results. Therefore, the compact object of HESS J1731-347 being an outlier in the distribution of observed pulsar masses may provide a unique insight to neutron star matter. However, one must take this estimation with caution as other studies have indicated several assumptions on the atmospheric composition and distance that may have led to these extreme numbers Alford and Halpern (2023). In particular, Alford and Halpern (2023) have pointed out that the results of Doroshenko et al. (2022) crucially depend on the following assumptions: i) the object has uniform-temperature carbon atmosphere, ii) the distance of this object is 2.5 kpc. If the distance is considered to be 3.2 kpc, the inferred mass would be much higher $\sim 1.4M_{\odot}$. They also have shown that the XMM-Newton spectra of this object is a poor fit with the assumed atmosphere model. As a consequence, further analyses are required to examine validity of the assumptions to fit the data. It is also important to mention that some latest supernova simulations are not able to find a remnant less than $1.19M_{\odot}$ Müller et al. (2025). Even though the results remain controversial, there has been a string of theoretical studies that investigates the nature of that object and its implications. For example, Horvath et al. (2023) have performed a minimal consistency check for the compact object of HESS J1731-347 within the existing theoretical models of compact stars. Zhang et al. (2025) have proposed a formation scenario for low mass neutron stars.
@@ II EOS of NS Matter
In this section, we provide a brief overview of our approach to model the supranuclear matter inside the NS. We have used the hybrid+PP EOS parametrization that has been developed by Biswas et al. (2021a) and used to provide multimessenger analyses of NS properties Biswas et al. (2021b); Biswas (2021). It has also been used to create a framework to infer Hubble parameter directly from GW signals from the future BNS mergers Ghosh et al. (2022, 2024). The core EOS consists of two main parts: a) nuclear physics informed expansion-based near saturation, b) agnostic polytropic at high densities. Around $\rho_{0}$, we have used an EOS based on Taylor’s series expansion of the isoscalar and isovector components of the energy per nucleon.
@@
where $\chi\equiv(\rho-\rho_{0})/3\rho_{0}$ expresses the deviation from $\rho_{0}$. We keep the terms upto second order as the higher order terms do not contribute significantly to the EOS around $\rho_{0}$. We also fix the saturation density, $\rho_{0}=0.16$ fm-3 and binding energy per nucleon of SNM at saturation, $e_{0}(\rho_{0})=-15.9$ MeV Brown and Schwenk (2014); Margueron et al. (2018); Tsang et al. (2019). Therefore the free parameters in this model are the incompressibility ($K_{0}$), nuclear symmetry energy ($e_{\rm{sym}}$) and its slope parameter ($L$), and the isovector incompressibility ($K_{\rm{sym}}$) at saturation. We consider the previous theoretical and experimental works to estimate these quantities Oertel et al. (2017); Dutra et al. (2012, 2014) and take a unifrom prior over a large range of value to incorporate their uncertainties. For the cold NS matter, the additional conditions of $\beta$-equilibrium and charge neutrality are also imposed. Then, at higher densities we have used an agnostic EOS with piecewise polytrope (PP) parametrization Read et al. (2009). In particular, we have used an arrangement with a three-segment polytrope ($\Gamma_{1},\Gamma_{2},\Gamma_{3}$ being the polytropic indices) starting at $1.25\rho_{0}$. The next two stitching points are $1.8$ and $3.6\rho_{0}$, respectively. Hence, we have a total of seven parameters, $\mathcal{E}=\{K_{0},e_{\rm{sym}},L,K_{\rm{sym}},\Gamma_{1},\Gamma_{2},\Gamma_{3}\}$, representing the EOS of NS matter. For the crust EOS, we have used the standard Baym-Pethick-Sutherland EOS table Baym et al. (1971).
@@ III Methodology
In this section, we describe briefly the inference methodology used in this work. We have used the tidal deformability posteriors from two BNS merger events GW1701817, and GW190425 Aasi et al. (2015); Acernese et al. (2015); Abbott et al. (2019, 2020). We have used the mass-radius posteriors of PSR J0030+0451 and PSR J0740+6620 Riley et al. (2019, 2021b); Miller et al. (2019b, 2021b). For the HESS J1731-347, we have used the M-R posterior Doroshenko et al. (2022) assuming the uniformly emitting carbon atmosphere provided by Doroshenko et al. (2022). In particular, we have used the data file, “xray_only_carbatm.txt” that includes only X-ray data using single temperature carbon atmosphere model and Gaia parallax priors. We have not used the other dataset that includes EOS priors from previous works as that would create an inconsistency with our own EOS model used in this work. The posterior of the EOS parameters can be expressed as,
@@ Acknowledgements
This work has been partially supported by the Fonds de la Recherche Scientifique-FNRS, Belgium, under grant No. 4.4501.19. PC is supported by European Union’s HORIZON MSCA-2022-PF-01-01 Programme under Grant Agreement No. 101109652, project ProMatEx-NS. BB acknowledges the support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2121 “Quantum Universe” – 390833306 and the Alexander von Humboldt Foundation through a Humboldt Research Fellowship for Postdoctoral Researchers.

### v3

# Compact object of HESS J1731-347 and its implication on neutron star matter

@@ I Introduction
Recent observation of a compact object inside the supernova remnant HESS J1731-347 has provided an estimation of its mass and radius to be $M=0.77^{+0.20}_{-0.17}M_{\odot}$ and $R=10.4^{+0.86}_{-0.78}$ km at $1\sigma$ credible levels, respectively Doroshenko et al. (2022). Understandably, this observation has generated a lot of attention from the nuclear physics community working on the dense matter physics beyond the normal nuclear matter density Brodie and Haber (2023); Huang et al. (2024); Li and Sedrakian (2023); Kubis et al. (2023); Koehn et al. (2025); Miao et al. (2024). Different possibilities have been explored to characterize this object, such as, a strange quark or a hybrid star Di Clemente et al. (2024); Horvath et al. (2023); Oikonomou and Moustakidis (2023); Rather et al. (2023); Yuan and Li (2024); Laskos-Patkos et al. (2024); Gao et al. (2024); Mariani et al. (2024), or a dark matter admixed compact star Sagun et al. (2023); Routaray et al. (2023). A typical neutron star (NS) usually has a mass within the range of $\sim 1-2M_{\odot}$. NSs being extremely dense compact objects have profound effects on our understanding of dense nuclear matter higher than nuclear saturation density ($\rho_{0}$) Glendenning (1997). The equation of state (EOS) of highly dense matter is connected to the observations of the masses and radii of the compact objects through the solutions of the Tolman-Oppenheimer-Volkoff equations. Hence, each new source with observed mass and radius allows us to test and constrain our models of nuclear physics and increase precision over the previous results. Therefore, the compact object of HESS J1731-347 being an outlier in the distribution of observed pulsar masses may provide a unique insight to neutron star matter. However, one must take this estimation with caution as other studies have indicated several assumptions on the atmospheric composition and distance that may have led to these extreme numbers Alford and Halpern (2023). In particular, Alford and Halpern (2023) have pointed out that the results of Doroshenko et al. (2022) crucially depend on the following assumptions: i) the object has uniform-temperature carbon atmosphere, ii) the distance of this object is 2.5 kpc. If the distance is considered to be 3.2 kpc, the inferred mass would be much higher $\sim 1.4M_{\odot}$. They also have shown that the XMM-Newton spectra of this object is a poor fit with the assumed atmosphere model. As a consequence, further analyses are required to examine validity of the assumptions to fit the data. It is also important to mention that some latest supernova simulations are not able to find a remnant less than $1.19M_{\odot}$ Müller et al. (2025). Even though the results remain controversial, there has been a string of theoretical studies that investigates the nature of that object and its implications. For example, Horvath et al. (2023) have performed a minimal consistency check for the compact object of HESS J1731-347 within the existing theoretical models of compact stars. Zhang et al. (2025) have proposed a formation scenario for low mass neutron stars.
@@ II EOS of NS Matter
In this section, we provide a brief overview of our approach to model the supranuclear matter inside the NS. We have used the hybrid nuclear + PP EOS parametrization that has been developed by Biswas et al. (2021a) and used to provide multimessenger analyses of NS properties Biswas et al. (2021b); Biswas (2021). It has also been used to create a framework to infer Hubble parameter directly from GW signals from the future binary neutron star BNS mergers Ghosh et al. (2022, 2024). The core EOS consists of two main parts: a) nuclear physics informed expansion-based near saturation, b) agnostic polytropic at high densities. Around $\rho_{0}$, we have used an EOS based on Taylor’s series expansion of the isoscalar and isovector components of the energy per nucleon.
@@
where $\chi\equiv(\rho-\rho_{0})/3\rho_{0}$ expresses the deviation from $\rho_{0}$. We keep the terms upto second order as the higher order terms do not contribute significantly to the EOS around $\rho_{0}$. We also fix the saturation density, $\rho_{0}=0.16$ fm-3 and binding energy per nucleon of SNM at saturation, $e_{0}(\rho_{0})=-15.9$ MeV Brown and Schwenk (2014); Margueron et al. (2018); Tsang et al. (2019). Therefore the free parameters in this model are the incompressibility ($K_{0}$), nuclear symmetry energy ($e_{\rm{sym}}$) and its slope parameter ($L$), and the isovector incompressibility ($K_{\rm{sym}}$) at saturation. We consider the previous theoretical and experimental works to estimate these quantities Oertel et al. (2017); Dutra et al. (2012, 2014) and take a unifrom prior over a large range of value to incorporate their uncertainties [see Table 1]. For the cold NS matter, the additional conditions of $\beta$-equilibrium and charge neutrality are also imposed. Then, at higher densities we have used an agnostic EOS with piecewise polytrope (PP) parametrization Read et al. (2009). In particular, we have used an arrangement with a three-segment polytrope ($\Gamma_{1},\Gamma_{2},\Gamma_{3}$ being the polytropic indices) starting at $1.25\rho_{0}$. The next two stitching points are $1.8$ and $3.6\rho_{0}$, respectively. Hence, we have a total of seven parameters, $\mathcal{E}=\{K_{0},e_{\rm{sym}},L,K_{\rm{sym}},\Gamma_{1},\Gamma_{2},\Gamma_{3}\}$, representing the EOS of NS matter. For the crust EOS, we have used the standard Baym-Pethick-Sutherland EOS table Baym et al. (1971).
@@ III Methodology
In this section, we describe briefly the inference methodology used in this work. We have used the tidal deformability posteriors from two BNS merger events GW1701817, and GW190425 Aasi et al. (2015); Acernese et al. (2015); Abbott et al. (2019, 2020). We have used the mass-radius posteriors of PSR J0030+0451 and PSR J0740+6620 Riley et al. (2019, 2021b); Miller et al. (2019b, 2021b). For the HESS J1731-347, we have used the mass-radius (M-R) posterior Doroshenko et al. (2022) assuming the uniformly emitting carbon atmosphere provided by Doroshenko et al. (2022). In particular, we have used the data file, “xray_only_carbatm.txt” that includes only X-ray data using single temperature carbon atmosphere model and Gaia parallax priors. We have not used the other dataset that includes EOS priors from previous works as that would create an inconsistency with our own EOS model used in this work. The posterior of the EOS parameters can be expressed as,
@@ Acknowledgements
This work has been partially supported by the Fonds de la Recherche Scientifique-FNRS, Belgium, under grant No. 4.4501.19. PC is supported by European Union’s HORIZON MSCA-2022-PF-01-01 Programme under Grant Agreement No. 101109652, project ProMatEx-NS. BB acknowledges the support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2121 “Quantum Universe” – 390833306 and the Alexander von Humboldt Foundation through a Humboldt Research Fellowship for Postdoctoral Researchers. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 101034371